When the play button is pressed, a CD accelerates uniformly from rest to 450 rev/min in 3.0 revolutions. If the CD has a radius of 7.0 cm and a mass of 15 g , what is the torque exerted on it? Express your answer using two significant figures.
Answer #1

Consider CD as a disk,

Moment of inertia is,

I = Mr2/2

= (0.015 kg)(0.07 m)2/2

= 3.675x10-5 kgm2

now,

ωf2 = ωi2 + 2αθ

[(450x2π)/(60 s)]2 = 02 + 2α(3x2π)

or, α = 58.9 rad/s2

torque is,

τ = Iα

= (3.675x10-5 kgm2)(58.9 rad/s2)

= 2.2x10-5 N.m

***********************************************************************************************
Check the answer and let me know in the comments immediately if you find something wrong or missing... I will rectify the mistakes asap if any